OpenStudy (anonymous):

Solving equations: 3x^2+6x+2=0

OpenStudy (zzr0ck3r):

\[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] a=3,b=6,c=2

OpenStudy (anonymous):

Yay I loveeee these -.- thanks for reminding me of the formula. I forgot it in a month of freedom .

OpenStudy (zzr0ck3r):

lol

OpenStudy (anonymous):

So the answer is (-3plus or minus the square root of 12)/6

OpenStudy (zzr0ck3r):

yep